Are the following lines parallel, perpendicular, or neither?
x-6y=12
6x+y=6

Answer by oscargut(2103)

x-6y=12
6x+y=6
first we can find y in each equation then
y = (x-12)/6 = (1/6)x-2
y = -6x+6 so coeficients of x are 1/6 and -6 then (1/6)(-6)=-1 then
the lines are perpendicular.
(lines ax+b and cx+d are perpendicular if only if ac=-1)